OpenStudy (anonymous):

A sample of gas is in a cylinder of fixed volume. At 30.0 °C its internal pressure is 3200. torr. What will be its pressure at 12.0 ° C?

OpenStudy (aaronq):

from the ideal gas law: \[PV=nRT \] \[\frac{ P _{1} }{ T _{1} }=\frac{ P _{2} }{ T _{2} }\] plug in and solve

OpenStudy (anonymous):

3010. torr?

OpenStudy (aaronq):

not quite. what did you do?

OpenStudy (joemc):

I got 3010 torr, too.

OpenStudy (anonymous):

Yea, I'm pretty sure that's the answer to the equation.

OpenStudy (aaronq):

\[\frac{ 3200 torr }{ 30 C }=\frac{ P _{2} }{ 12C } \to P _{2}=\frac{ 3200 torr }{ 30 C }*12C \to P _{2}= 1280 torr\]

OpenStudy (joemc):

@aaronq, you have to use Kelvin.....

OpenStudy (joemc):

\[\frac{ 3200 torr }{303.15 K } = \frac{ P_2 }{ 285.15 K } = \frac{3200 torr }{303.15 K } \times 285.15 K = P_2\]

OpenStudy (aaronq):

haha wow. FAIR ENOUGH